# Structured World Modeling via Semantic Vector Quantization

### v1

@@ Abstract
Neural discrete representations are crucial components of modern neural networks. However, their main limitation is that the primary strategies such as VQ-VAE can only provide representations at the patch level. Therefore, one of the main goals of representation learning, acquiring structured, semantic, and compositional abstractions such as the color and shape of an object, remains elusive. In this paper, we present the first approach to semantic neural discrete representation learning. The proposed model, called Semantic Vector-Quantized Variational Autoencoder (SVQ), leverages recent advances in unsupervised object-centric learning to address this limitation. Specifically, we observe that a simple approach quantizing at the object level poses a significant challenge and propose constructing scene representations hierarchically, from low-level discrete concept schemas to object representations. Additionally, we suggest a novel method for structured semantic world modeling by training a prior over these representations, enabling the ability to generate images by sampling the semantic properties of the objects in the scene. In experiments on various 2D and 3D object-centric datasets, we find that our model achieves superior generation performance compared to non-semantic vector quantization methods such as VQ-VAE and previous object-centric generative models. Furthermore, we find that the semantic discrete representations can solve downstream scene understanding tasks that require reasoning about the properties of different objects in the scene.
@@ 1 Introduction
While there have been various findings regarding the purpose of the brain, it is fair to say that the human brain has at least two key functions. First, it constructs a good representation that captures the structure of the world through perception. Second, it imagines or generates various possibilities of the world. Similarly, AI systems that aim to be as generally capable as humans would also need to realize similar capabilities computationally. Building such a learning system that can both structurally recognize and generate has long been a desired vision in machine learning, from Helmholtz machines (Dayan et al., 1995; Wade, 2021) to Variational Autoencoders (Kingma & Welling, 2013; Rezende et al., 2014). Although there could be various approaches to achieving this, in this work, we focus on a specific type of modeling, which we call Generative Structured World Modeling, which satisfy the following desiderata.
@@
First, when it comes to representating a visual scene, it appears that we do not perceive the scene simply as a monolithic vector of features. Instead, we view it structurally and semantically, recognizing it as a composition of meaningful components such as objects and their attributes like shape, color, and position (Palmer, 1977; Singer, 2007; Spelke & Kinzler, 2007). Various works in AI, particularly object-centric approaches (Greff et al., 2020), have demonstrated that this structural decomposition facilitates relational reasoning (Wu et al., 2021; Yoon et al., 2023; Webb et al., 2023a; b) and out-of-distribution generalization (Dittadi et al., 2022; Yoon et al., 2023) due to improved compositional generalization. It has also been shown that a monolithic vector representation of a scene, such as VAE, fails in multi-object scenes (Wu et al., 2021; Dittadi et al., 2022; Yoon et al., 2023).
@@
Moreover, this structured and semantic understanding can be categorized and conceptualized discretely in an unsupervised way. Such an ability is critical for organizing and comprehending the complexity of the environment, e.g., via language, as well as for implementing modularity (Andreas et al., 2016) or symbolic reasoning (Lake et al., 2016). In AI, discrete representations are also useful to leverage powerful learning models like transformers. One of the most popular models for discrete representation learning in AI is VQ-VAE (van den Oord et al., 2017). It has been shown to be beneficial for image generation (Razavi et al., 2019; Esser et al., 2021) and probability density modeling (Van den Oord et al., 2016). However, VQ-VAE and its variants, such as dVAE (Ramesh et al., 2021; Singh et al., 2022a) and VQ-GAN (Esser et al., 2021), represent a scene as a grid of small patches, lacking the capability to capture the scene’s holistic structure and semantics.
@@
Besides, the ability to generate samples that adhere to the observed data distribution is foundational for endowing AI the capabilities to imagine and simulate, e.g., for planning (Mattar & Lengyel, 2022; Hafner et al., 2019). However, only a certain class of representation learning models supports this essential ability. While models like Slot Attention (Locatello et al., 2020) and SysBinder (Singh et al., 2023) offer structured, object-centric representations, in its original form it is unclear how to support density-based sampling. In contrast, VAE-based models, such as VAE and VQ-VAE, generally support this ability to sample from a prior distribution, but they either do not provide object-centric structures (VAE) or are limited to patch-based representations (VQ-VAE).
@@
In this work, we observe that, despite its significance, there is currently no method that simultaneously satisfies all of the mentioned criteria of Generative Structured World Modeling, as summarized in Table 1. To address this issue, we propose the Semantic Vector-Quantized (SVQ) Variational Autoencoder. Our model achieves discrete semantic decomposition by learning hierarchical, composable factors that closely align with the objects and the properties of objects in visual scenes. Similar to patch-based vector quantization methods, we can train an autoregressive prior to learn the distribution of the dataset. However, unlike VQ-VAE, we achieve this by learning the distribution of semantic discrete tokens, rather than patch tokens. As a result, the generation (or imagination) process is to compose semantic concepts such as objects and their attributes, rather than stitching a grid of patches.
@@
In our experiments, we demonstrate the following practical benefits of our method. First, we find that for multi-object scenes, SVQ is able to model the prior distribution better than patch-based methods, as measured by the quality of the samples generated. Second, we find that SVQ representations outperform patch-based VQ representations on downstream tasks that require knowledge of the different properties of the objects in the scene. We also find evidence that SVQ representations can generalize better to out-of-distribution tasks compared to patch-based VQ representations and SysBinder continuous representations. Lastly, we show that despite introducing a discrete bottleneck, SVQ can work on the challenging CLEVRTex (Karazija et al., 2021) dataset, one of the most complex datasets used in recent unsupervised object-centric representation learning models.
@@
Our contributions are as follows: First, we introduce SVQ, the first model to obtain semantic neural discrete representations without any supervision about the underlying factors in the scene. Second, by training a prior over these discrete representations, we are able to obtain an object-centric density model, capable of capturing the underlying data distribution and generating new samples. Third, we evaluate our model on several 2D and 3D datasets including the challenging CLEVRTex dataset, showing superior downstream task performance and image generation quality.
Table 1: Desiderata for Generative Structured World Modeling and Related Models
VAE VQ-VAE Slot Attention SysBinder SVQ (Ours) Semantic Decomposition Factor Object Object & Factor Object & Factor Discrete Sampling
@@ 3.1 Semantic Vector Quantization
Given a slot attention encoder that can obtain a set of representations of the objects in a scene, one may think of a hypothetical method, applying vector quantization to the slot representation itself to obtain a set of semantic discrete representations (Figure 1b). While these representations would indeed correspond to the different objects in a scene, this scheme would essentially require one codebook entry per possible object configuration and would be insufficient for anything beyond trivially simple scenes.
For example, consider a simple scene containing a single object in a fixed position that only varies by color and shape. Assume there are $c$ possible colors and $s$ possible shapes for the object. With slot-level quantization, in order to represent all the potential objects, the codebook would require at least $c\times s$ entries. This is because each slot representation is a single entangled representation so each combination of factors needs to be represented by a separate code. If instead, we were able to disentangle the object-level representations into factor-level representations—representations that align with the underlying latent factors of variation of each object—we would be able to describe the potentially large combinatorial space of each object with a much small number of discrete factors. In the above example, if we had a fully disentangled representation of the color and the shape, we would be able to represent all possible scenes with $c+s$ codes (Figure 1c). See Appendix A.2 for further discussion.
@@
Figure 2: (a) Overall architecture of SVQ. We maintain $M$ learned codebooks and split each slot into $M$ blocks. At the end of each Slot Attention iteration, we apply vector quantization to each block representation to obtain a set of discrete codes for each slot. Each block ends up specializing in different underlying factors of variation for the objects in the scene. (b) The Autoregressive Semantic Prior. After training the model, we freeze SVQ and train and autoregressive prior over the discrete latent codes. Sampling from this prior allows us to generate an image one object at a time, based on their properties.
3.2 Autoregressive Semantic Prior
Given these semantic discrete codes representing the different objects in the scene, we can now freeze the SVQ and train an autoregressive prior $p(\textbf{z}_{q})$ over these codes to obtain a generative model of the underlying data that captures the structure and semantics of the data. We can then sample from this prior to obtain codes for new scenes and use these codes in the SVQ decoder to generate new images. Compared to patch-based VQ methods that generate tokens that correspond to a spatially local region of an image, this Autoregressive Semantic Prior (ASP) generates an image one object at a time, based on their properties (Figure 2b).
@@ 5 Experiments
Baselines. We compare our model with several patch-based quantization methods: VQ-VAE (van den Oord et al., 2017) with a PixelCNN (Van den Oord et al., 2016) prior, and dVAE (Ramesh et al., 2021; Singh et al., 2022a) with a transformer decoder prior. For the dVAE baseline, we use the dVAE weights that are trained along with the SVQ. This provides a more direct ablation comparing the ASP of SVQ with the patch-based transformer decoder prior since the dVAE decoder is shared across these models and will not contribute to differences in image quality. We also compare with GENESIS-v2 (Engelcke et al., 2022), a continuous latent object-centric model with an autoregressive prior that can also generate samples.
5.1 Generating Samples with the Autoregressive Semantic Prior
@@
We see that for the simplest dataset with 3 objects and no background, SVQ achieves the lowest FID of the models and comparable generation accuracy to dVAE, generating about 75% of the scenes correctly. This setting may be simple enough that dVAE with a transformer prior can capture the structure of the scene even with a patch-based discrete latent. As the scene complexity increases with more objects and textured background, SVQ starts to outperform the baselines in terms of generation accuracy. Inspecting the qualitative results, we see that in the dataset with the background, VQ-VAE and dVAE start generating occasional blurry objects, whereas SVQ maintains clean-looking objects that match the ground truth dataset. This may be because SVQ can segment the background into its own slot and factor the texture into a discrete latent, cleanly separating the representation of the objects from the background. The patch-based methods, however, may have a harder time separating the foreground from the background resulting in messier generations. Interestingly, despite the blurry shapes, VQ-VAE achieves the lowest FID score on the 2D Sprites dataset with background. We hypothesize this may be because the model spends more capacity modeling the background correctly instead of the foreground, which may produce a better FID score, but not necessarily better generation accuracy. This is confirmed by the low generation accuracy of the VQ-VAE model this dataset, only generating 19.5% of the scenes correctly.
@@
FID $\downarrow$ Generation Accuracy (in %) $\uparrow$ Dataset VQ-VAE dVAE SVQ (ours) VQ-VAE dVAE SVQ (ours) 2D Sprites (3 obj) 14.81 7.26 6.61 28.91 75.78 75.00 2D Sprites (4 obj) 26.35 19.15 17.93 21.88 62.50 66.41 2D Sprites w/ bg (4 obj) 58.14 66.08 58.50 19.53 30.47 42.19
@@ 5.1.2 CLEVR
In Figure 4, we show sample generations after training the models on the CLEVR-Easy, CLEVR-Hard, and CLEVR-Tex datasets. We report the Frechet Inception Distance (FID) in Table 3. We find that compared to the other models, GENESIS-v2 generates very blurry images and completely fails on CLEVR-Tex, resulting in a high FID. While VQ-VAE produces sharper images, several of the generated shapes are malformed or have mixed colors. The dVAE-generated images look closer to the ground truth dataset, but still have some errors such as overlapping objects (first image) and generating scenes with more objects than seen in the training set (third image). SVQ has the lowest FID for all of these datasets and the generated images look very close to the ground truth dataset, indicating the usefulness of the ASP for generating these multi-object scenes.
In Appendix A.1, we show additional analysis of the learned codebook on the CLEVR-Easy dataset.
@@
FID $\downarrow$ Dataset GENESIS-v2 VQ-VAE dVAE SVQ (ours) CLEVR-Easy 115.56 57.06 40.30 32.50 CLEVR-Hard 93.01 73.33 65.89 43.12 CLEVR-Tex 225.08 178.59 112.80 84.52
@@ 5.2.1 Odd-One-Out
We first evaluate on a downstream supervised learning task on the 2D Sprites dataset. We modify the dataset by first dividing each image into four quadrants and ensuring exactly one object will be in each quadrant. As in our previous experiments, one object has a single property that is unique from the other objects. The goal of the task is to identify the quadrant of the odd-one-out object. We first pretrain the baseline models on a dataset containing all 12 possible shapes and 7 possible colors. Then, we freeze the underlying model and train a downstream model on top of the learned representations with the supervised objective. The downstream model is trained on a dataset that only contains 9 possible shapes and 4 possible colors. We then evaluate on both the in-distribution (ID) dataset and an out-of-distribution (OOD) dataset that consists of the remaining 3 shapes and 3 colors. In addition to dVAE and VQ-VAE, we use SysBinder as a baseline for this task, to compare its continuous representation with SVQ’s discrete representation. For the latent representation of SVQ, we include variants that use the codebook indices (SVQ Indices) and the codebook prototype vectors (SVQ Codebook).
@@
Steps to 98% ($\downarrow$) OOD Acc. % ($\uparrow$) dVAE Discrete 37,000 26.7 dVAE Continuous 32,000 29.5 VQ-VAE Indices 77,000 24.0 VQ-VAE Codebook 54,500 55.6 SysBinder 27,000 67.6 SVQ Indices 77,000 46.8 SVQ Codebook 27,000 99.1
ID Acc. % ($\uparrow$) OOD Acc. % ($\uparrow$) dVAE Discrete 27.52 19.87 dVAE Continuous 24.51 20.51 VQ-VAE Indices 24.53 17.74 VQ-VAE Codebook 23.73 18.80 SysBinder 79.60 70.09 SVQ Indices 68.21 64.53 SVQ Codebook 75.86 71.15
Table 5 shows the results of our experiments. Since all models can solve the task when evaluated on the ID dataset, we report the number of steps to reach 98% accuracy on the validation dataset. We find that SysBinder and SVQ Codebook learn the quickest in the ID setting. For the OOD setting, we find that dVAE and VQ-VAE fail completely, not performing better than randomly guessing, showing that the patch-based discrete latent is insufficient for OOD generalization in this task. SysBinder can partially solve the task in the OOD setting, while the SVQ Codebook seems to be able to solve the task, achieving 99% accuracy. This indicates that the compact latent space offered by the discrete code provides better out-of-distribution generalization abilities for this particular task. One possible explanation for this is that since this is an odd-one-out task, the downstream network needs to do comparisons between the properties of the objects and this may be easier to do with SVQ’s codebook vectors that are fixed. SysBinder’s continuous latents, on the other hand, offer greater variations for the same concept. This increases the potential for the downstream network to learn spurious correlations in the data, which can negatively impact OOD performance. SVQ Indices is also only able to partially solve the task. This makes sense because in the out-of-distribution case, the model does not have any way of knowing two codebook indices are for the same property value (e.g. if two codebook vectors both correspond to the color blue). Since SVQ Codebook uses the prototype vectors, it does not have this problem because the similarity can be determined by the vector representation.

### v2

# Neural Language of Thought Models

@@ Abstract
The Language of Thought Hypothesis suggests that human cognition operates on a structured, language-like system of mental representations. While neural language models can naturally benefit from the compositional structure inherently and explicitly expressed in language data, learning such representations from non-linguistic general observations, like images, remains a challenge. In this work, we introduce the Neural Language of Thought Model (NLoTM), a novel approach for unsupervised learning of LoTH-inspired representation and generation. NLoTM comprises two key components: (1) the Semantic Vector-Quantized Variational Autoencoder, which learns hierarchical, composable discrete representations aligned with objects and their properties, and (2) the Autoregressive LoT Prior, an autoregressive transformer that learns to generate semantic concept tokens compositionally, capturing the underlying data distribution. We evaluate NLoTM on several 2D and 3D image datasets, demonstrating superior performance in downstream tasks, out-of-distribution generalization, and image generation quality compared to patch-based VQ-VAE and continuous object-centric representations. Our work presents a significant step towards creating neural networks exhibiting more human-like understanding by developing LoT-like representations and offers insights into the intersection of cognitive science and machine learning.
@@ 1 Introduction
The Language of Thought Hypothesis (LoTH) (Fodor et al., 1975) suggests that human cognition is based on a structured, language-like system of mental representations, often referred to as “Mentalese”. Mentalese comprises word-like units that form sentence-like structures, which convey meaning. The meaning of these mental “sentences” is systematically determined by the meanings of their constituent “words” and their specific arrangement. From a computational viewpoint, while neural language models (Bengio et al., 2000; Brown et al., 2020; Bommasani et al., 2021) can benefit from the compositional and symbolic structure inherently expressed in the language data they are trained on, it remains unclear how we can learn such LoT-like structure from non-linguistic general observations, such as images, videos, and audio signals. The significance of this ability is further highlighted by the fact that infants learn these structures from observing objects and events before they acquire language skills (Spelke, 2022).
How can we create neural networks that learn to develop such language of thought representations in an unsupervised way? To address this, we outline the following three properties as the desired characteristics of a neural language of thought model.
@@
First, when perceiving a visual scene, humans do not simply represent it as a monolithic vector of features. Instead, we view the scene structurally and semantically, recognizing it as a composition of meaningful components such as objects and their attributes, including shape, color, and position (Palmer, 1977; Singer, 2007; Spelke & Kinzler, 2007). Our observation here is that in line with the LoTH, these visual attributes can be likened to words, objects to sentences, and the scene to a paragraph. Recent works, particularly those focused on object-centric representations (Greff et al., 2020), have demonstrated that this structural decomposition facilitates the benefits associated with the LoTH such as relational reasoning (Wu et al., 2021; Yoon et al., 2023; Webb et al., 2023a; b) and out-of-distribution generalization (Dittadi et al., 2022; Yoon et al., 2023) due to increased compositional generalization.
@@
Moreover, these structured and semantic representations can be categorized and conceptualized, resulting in symbol-like discrete concept abstraction. Such an ability is critical for organizing and comprehending the complexity of the environment, e.g., via language, as well as for implementing modularity (Andreas et al., 2016) or symbolic reasoning (Lake et al., 2016). Such discrete representations are also useful to leverage powerful generative models like autoregressive transformers (Vaswani et al., 2017). One of the most popular models for discrete representation learning is VQ-VAE (van den Oord et al., 2017). It has been shown to be beneficial for image generation (Razavi et al., 2019; Esser et al., 2021) and probability density modeling (Van den Oord et al., 2016). However, VQ-VAE and its variants, such as dVAE (Ramesh et al., 2021; Singh et al., 2022a) and VQ-GAN (Esser et al., 2021), represent a scene as a grid of small patches, lacking the capability to capture the scene’s holistic structure and semantics.
@@
Besides, the ability to compositionally and probabilistically generate samples that adhere to the distribution of prior beliefs, constructed from observation data, is crucial for endowing AI with the capabilities to imagine and simulate. These abilities are essential for tasks such as planning (Mattar & Lengyel, 2022; Hafner et al., 2019) and reasoning. However, this ability, related to probabilistic Language of Thought (PLoT) (Goodman et al., 2015) and productivity, is supported by only a certain class of representation learning models. While models like Slot Attention (Locatello et al., 2020) and SysBinder (Singh et al., 2023) offer structured, object-centric compositional representations, in their original form it is unclear how they support density-based sampling. In contrast, VAE-based models support this ability to sample from a prior distribution. However, they either do not provide object-centric structures or are limited to patch-based discrete abstractions (VQ-VAE).
Table 1: Desiderata for Neural Language of Thought Models and Related Models
VAE VQ-VAE Slot Attention SysBinder NLoTM (Ours) Compositionality (Semantic Scene Decomposition) Factor Object Object & Factor Object & Factor Symbolic (Discrete Concept Abstraction) (Patch Concept) (Semantic Concept) Productivity (Probabilistic Compositional Generation) (Patch Stitching) (Semantic Composition)
@@
In this work, we present the Neural Language of Thought Model (NLoTM), the first model that satisfies the aforementioned criteria summarized in Table 1. NLoTM comprises two novel components: the Semantic Vector-Quantized (SVQ) Variational Autoencoder and the Autoregressive LoT Prior (ALP). SVQ achieves discrete semantic decomposition of a scene by learning hierarchical, composable factors that closely align with the objects and the properties of objects in visual scenes, similar to the role of words and sentences in LoTH. ALP, analogous to language models trained on text tokens, is an autoregressive transformer trained to learn a probabilistic generation of semantic concept tokens. However, unlike VQ-VAE which stitches a grid of patches, ALP composes semantic concepts such as objects and their attributes to represent a scene.
@@
In our experiments, we demonstrate the following practical benefits of our method. First, we find that for multi-object scenes, NLoTM is able to model the prior distribution better than patch-based methods, as measured by the quality of the samples generated. Second, we find that NLoTM representations outperform patch-based VQ representations on downstream tasks that require knowledge of the different properties of the objects in the scene. We also find evidence that NLoTM representations can generalize better to out-of-distribution tasks compared to patch-based VQ representations and SysBinder continuous representations. Lastly, we show that despite introducing a discrete bottleneck, NLoTM can work on the challenging CLEVRTex (Karazija et al., 2021) dataset, one of the most complex datasets used in recent unsupervised object-centric representation learning models.
@@
Our contributions are as follows: First, we introduce NLoTM, the first neural network model implementing the LoTH for unstructured observations. Second, we propose the SVQ model to obtain object-centric semantic neural discrete representations. Third, we propose the ALP to model the probabilistic prior capable of capturing the underlying data distribution and compositional generation of new samples. Lastly, we evaluate our model on several 2D and 3D datasets including the challenging CLEVRTex dataset, showing superior downstream task performance and image generation quality.
@@ 3.1 Semantic Vector Quantization
Given a slot attention encoder that can obtain a set of representations of the objects in a scene, one may think of a hypothetical method, applying vector quantization to the slot representation itself to obtain a set of semantic discrete representations (Figure 1b). While these representations would indeed correspond to the different objects in a scene, this scheme would require one codebook entry per possible object configuration and be insufficient for anything beyond trivially simple scenes.
For example, consider a simple scene containing a single object in a fixed position that only varies by color and shape. Assume there are $c$ possible colors and $s$ possible shapes for the object. With slot-level quantization, in order to represent all the potential objects, the codebook would require at least $c\times s$ entries. This is because each slot representation is a single entangled representation, so each combination of factors needs to be represented by a separate code. If, instead, we were able to disentangle the object-level representations into factor-level representations—representations that align with the underlying latent factors of variation of each object—we could describe the potentially large combinatorial space of each object with a much smaller number of discrete factors. In the above example, if we had a fully disentangled representation of the color and the shape, we could represent all possible scenes with $c+s$ codes (Figure 1c). See Appendix B.2 for further discussion.
@@
Figure 2: Overall architecture of NLoTM. (a) The Semantic Vector-Quantized (SVQ) Variational Autoencoder. We maintain $M$ learned codebooks and split each slot into $M$ blocks. After each Slot Attention iteration, we apply vector quantization to each block representation to obtain a set of discrete codes for each slot. Each block ends up specializing to different underlying factors of variation for the objects in the scene. (b) The Autoregressive LoT Prior (ALP). We train an autoregressive prior over the discrete latent codes from SVQ. Sampling from this prior allows us to generate an image one object at a time, based on their properties.
3.2 Autoregressive Language of Thought Prior
Given these semantic discrete codes representing the different objects in the scene, we can now freeze the SVQ and train a prior $p(\textbf{z}_{q})$ over these codes to obtain a generative model of the underlying data that captures the structure and semantics of the data, analogous to language models that are trained over text tokens. We can sample from this prior, called Autoregressive LoT Prior (ALP), to obtain codes for new scenes and use these codes in the SVQ decoder to generate new images. Compared to patch-based VQ methods that generate patch tokens corresponding to a local region of an image, ALP generates a representation of a scene one object at a time, based on their properties (Figure 2b).
@@ 5 Experiments
Baselines. We compare our model with several patch-based quantization methods: VQ-VAE (van den Oord et al., 2017) with a PixelCNN (Van den Oord et al., 2016) prior, and dVAE (Ramesh et al., 2021; Singh et al., 2022a) with a transformer decoder prior. For the dVAE baseline, we use the dVAE weights that are trained along with the SVQ. This provides a more direct ablation comparing the ALP of NLoTM with the patch-based transformer decoder prior since the dVAE decoder is shared across these models and will not contribute to differences in image quality. We also compare with GENESIS-v2 (Engelcke et al., 2022), a continuous latent object-centric model with an autoregressive prior that can also generate samples.
5.1 Generating Samples with the Autoregressive LoT Prior
@@
We see that for the simplest dataset with 3 objects and no background, NLoTM achieves the lowest FID of the models and comparable generation accuracy to dVAE, generating about 75% of the scenes correctly. This setting may be simple enough that dVAE with a transformer prior can capture the structure of the scene even with a patch-based discrete latent. As the scene complexity increases with more objects and textured background, NLoTM starts to outperform the baselines in terms of generation accuracy. Inspecting the qualitative results, we see that in the dataset with the background, VQ-VAE and dVAE start generating occasional blurry objects, whereas NLoTM maintains clean-looking objects that match the ground truth dataset. This may be because NLoTM can segment the background into its own slot and factor the texture into a discrete latent, cleanly separating the representation of the objects from the background. The patch-based methods, however, may have a harder time separating the foreground from the background resulting in messier generations. Interestingly, despite the blurry shapes, VQ-VAE achieves the lowest FID score on the 2D Sprites dataset with background. We hypothesize this may be because the model spends more capacity modeling the background correctly instead of the foreground, which may produce a better FID score, but not necessarily better generation accuracy. This is confirmed by the low generation accuracy of the VQ-VAE model this dataset, only generating 19.5% of the scenes correctly.
@@
FID $\downarrow$ Generation Accuracy (in %) $\uparrow$ Dataset VQ-VAE dVAE NLoTM (ours) VQ-VAE dVAE NLoTM (ours) 2D Sprites (3 obj) 14.81 7.26 6.61 28.91 75.78 75.00 2D Sprites (4 obj) 26.35 19.15 17.93 21.88 62.50 66.41 2D Sprites w/ bg (4 obj) 58.14 66.08 58.50 19.53 30.47 42.19
@@ 5.1.2 CLEVR
In Figure 4, we show sample generations after training the models on the CLEVR-Easy, CLEVR-Hard, and CLEVR-Tex datasets. We report the Frechet Inception Distance (FID) in Table 3. We find that compared to the other models, GENESIS-v2 generates very blurry images and completely fails on CLEVR-Tex, resulting in a high FID. While VQ-VAE produces sharper images, several of the generated shapes are malformed or have mixed colors. The dVAE-generated images look closer to the ground truth dataset, but still have some errors such as overlapping objects (first image) and generating scenes with more objects than seen in the training set (third image). NLoTM has the lowest FID for all of these datasets and the generated images look very close to the ground truth dataset, indicating the usefulness of the ASP for generating these multi-object scenes.
In Appendix B.1, we show additional analysis of the learned codebook on the CLEVR-Easy dataset.
@@
FID $\downarrow$ Dataset GENESIS-v2 VQ-VAE dVAE NLoTM (ours) CLEVR-Easy 115.56 57.06 40.30 32.50 CLEVR-Hard 93.01 73.33 65.89 43.12 CLEVR-Tex 225.08 178.59 112.80 84.52
@@ 5.2.1 Odd-One-Out
We first evaluate on a downstream supervised learning task on the 2D Sprites dataset. We modify the dataset by first dividing each image into four quadrants and ensuring exactly one object will be in each quadrant. As in our previous experiments, one object has a single property that is unique from the other objects. The goal of the task is to identify the quadrant of the odd-one-out object. We first pretrain the baseline models on a dataset containing all 12 possible shapes and 7 possible colors. Then, we freeze the underlying model and train a downstream model on top of the learned representations with the supervised objective. The downstream model is trained on a dataset that only contains 9 possible shapes and 4 possible colors. We then evaluate on both the in-distribution (ID) dataset and an out-of-distribution (OOD) dataset that consists of the remaining 3 shapes and 3 colors. In addition to dVAE and VQ-VAE, we use SysBinder as a baseline for this task, to compare its continuous representation with NLoTM’s discrete representation. For the latent representation of NLoTM, we include variants that use the codebook indices in the SVQ (NLoTM Indices) and the codebook prototype vectors in the SVQ (NLoTM Codebook).
@@
Steps to 98% ($\downarrow$) OOD Acc. % ($\uparrow$) dVAE Discrete 37,000 26.7 dVAE Continuous 32,000 29.5 VQ-VAE Indices 77,000 24.0 VQ-VAE Codebook 54,500 55.6 SysBinder 27,000 67.6 NLoTM Indices 77,000 46.8 NLoTM Codebook 27,000 99.1
@@
ID Acc. % ($\uparrow$) OOD Acc. % ($\uparrow$) dVAE Discrete 27.52 19.87 dVAE Continuous 24.51 20.51 VQ-VAE Indices 24.53 17.74 VQ-VAE Codebook 23.73 18.80 SysBinder 79.60 70.09 NLoTM Indices 68.21 64.53 NLoTM Codebook 75.86 71.15
Table 5 shows the results of our experiments. Since all models can solve the task when evaluated on the ID dataset, we report the number of steps to reach 98% accuracy on the validation dataset. We find that SysBinder and NLoTM Codebook learn the quickest in the ID setting. For the OOD setting, we find that dVAE and VQ-VAE fail, not performing better than randomly guessing, showing that the patch-based discrete latent is insufficient for OOD generalization in this task. SysBinder can partially solve the task in the OOD setting, while the NLoTM Codebook seems to be able to solve the task, achieving 99% accuracy. This indicates that the compact latent space offered by the discrete code provides better OOD generalization abilities for this particular task. One possible explanation for this is that since this is an odd-one-out task, the downstream network needs to do comparisons between the properties of the objects and this may be easier to do with SVQ’s codebook vectors that are fixed. SysBinder’s continuous latents, on the other hand, offer greater variations for the same concept. This increases the potential for the downstream network to learn spurious correlations in the data, which can negatively impact OOD performance. NLoTM Indices is also only able to partially solve the task. This makes sense because in the out-of-distribution case, the model does not have any way of knowing two codebook indices are for the same property value (e.g. if two codebook vectors both correspond to the color blue). Since NLoTM Codebook uses the prototype vectors, it does not have this problem because the similarity can be determined by the vector representation.